the lenghs of a rectangleis twice the the width. the perimeter is 84 cm more than the width. make a digram and find the rectangle's dimensions.
Answer by ankor@dixie-net.com(22740)

:
the lengths of a rectangle is twice the the width.
L = 2W
the perimeter is 84 cm more than the width.
2L + 2W = W + 84
:
Substitute 2W for L (from the 1st equation)
2(2W) + 2W = W + 84
4W + 2W - W = 84
5W = 84
W = 84/5
W = 16.8 cm
:
L = 2 * 16.8 = 33.6 cm
:
A rectangle 33.6 by 16.8
:
Check it using the perimeter:
2(33.6) + 2(16.8) = 100.8 which is 84 cm more than (16.8). the width
:
:
I'll let you make a digram.